# Blockchain Economic Denial of Sustainability Attack: Exploiting Latency Optimization in Ethereum Transaction Forwarding

### v1

@@ 4.3 Network waste
To estimate how one invalid transaction amplifies in the network, we first derive 1) the number of bytes a regular node $i$ receives based on the number of peer connections (see blue arrows directed toward each regular node in Figure 1), 2) the number of bytes the entire network wastes based on the distributions of active connections on each peer (all the blue arrows). This calculation only consider the public nodes in a P2P network, but not the users (MEV bots and block miners/builders) who 1) privately connect to txpool providers or 2) listen to transactions through RPC endpoints outside the P2P network. The actual amplification factor would be the combination of 1) the amount of traffic that goes to users who listen to modified nodes and 2) TAF. We indeed notice that some modified nodes shut down the outbound connections to reduce outbound costs. Yet, they still incur the cost of propagating invalid transactions to users.
@@
As we already determine the waste for each regular node with a varying number of active connections, we proceed to calculate the total waste in the network by considering the distribution of peer connections, $g(x)$, which we empirically estimate in §6.1. There are $(1-\gamma)\cdot N=N_{\mathcal{R}}$ regular nodes where $N$ is the total number of nodes in the network. We multiply the number of regular nodes in the network, the distributions of peer connections $g(x)$, and the amount of waste for each node ($f(x)$) to get the overall network waste:
[…]
where $\int_{0}^{\infty}f(x)g(x)$ represents the expected waste for one average regular node. In practice, we integrate until 1,000 connections (to ensure convergence and exclude an impractically large number of connections) using Scipy. We finally divide this by the transaction size ($a$) to get TAF. An adversary could set up modified nodes on their own to increase $\gamma$, but these nodes would also be attacked, making this strategy unproductive.
@@ 5 Empirical examples of attack
This section identifies the similar attack instances seen in the wild, and characterizes the size, intensity, cost, and strategies, and actors (the attacker – sender address, the victim – centralized services) involved in those attacks. While we cannot definitively pinpoint the motivations of those attackers, those attacks are executed in a manner consistent with our EDoS attack (see the discussion in Appendix C).
@@ 5.1 Victims
Figure 5 (right) is a co-occurrence matrix where each cell indicates the number of dropped transactions common to both services (the $x$- and $y$-axis) normalized by each day. The diagonal is the number of dropped transactions to each source—e.g., bloXroute has around 60,000 new dropped transactions per day. bloXroute, Chainbound, and Eden collectively accept a significantly high number of dropped transactions that are not seen in any of the other txpools, which suggests that they may have employed different propagation strategies from others. Those results hypothesize that a part of the latency reduction might also come from transaction validation process, which motivates our simulations in §8.1.
Figure 5: Comparing the centralized services’ txpools (left: accuracy-latency plot, right: dropped transactions between sources).
@@ 5.3 Characterizing attack behavior
We first check the three types of invalid transactions discussed in §4.1: 1) insufficient balance, 2) past nonce, and 3) duplicate transaction. To check the balance and nonce of the relevant accounts at the time of transaction submission, we use our archive node to fetch the state of the blockchain at the time of the timestamp. Given that there is a network delay between when the transaction enters a P2P network and when Flashbots records the timestamp, we look at the state of the account two blocks before the observation timestamp (i.e., at least 12 seconds in the past). Of 2,591 attack cases, we confirm that 536 instances (from 224 unique addresses) come from accounts that do not have a sufficient balance. Two services—bloXroute and Eden—have included those transactions during our observation window. In particular, some addresses do not have any transactions (send or receive) in their history whatsoever (i.e., zero Ether balance). In addition, three services (Chainbound, Eden, bloXroute) have accepted transactions that were using a past nonce (and thus, are invalid) during our measurement interval. Just like we identify invalid transactions from accounts with insufficient balance, we retrieve the last nonce of the corresponding address and identify 62 attacks from 27 unique addresses that use previously used nonce.
@@ 6.1 Peer connection distribution: $g(x)$
We formalize our method for estimating the number of active connections for each peer, $x_{i}$, and empirically computing its distribution, $g(x)$. Although studies on blockchain P2P network topology exist, none of the approaches directly translate to our context. The key novelty in our approach is to estimate active peer connections on the Ethereum mainnet without submitting test transactions (i.e., mediating ethical concerns and research costs). We use a 5-month dataset, encompassing 2.5 billion messages.
As outlined in §4, a unmodified Geth node $i$ has a number of peer connections $x_{i}$ and uniformly randomly chooses $\sqrt{x_{i}}$ nodes to broadcast (0x02) transactions, while sending announces (0x08) to the rest. Our node receives a 0x02 message with probability $\theta_{i}=\frac{\sqrt{x_{i}}}{x_{i}}$, which means the data distribution follows a binomial distribution with parameters $\theta_{i}$ and $m$, where $m$ is the number of messages we receive from each peer. Specifically, $m=m_{2}+m_{8}$ where $m_{2}$ and $m_{8}$ are the number of 0x02 and 0x08 messages sent by peer $i$, respectively. The probability (likelihood function) of the binomial distribution is defined as
@@
We derive an unbiased MLE (Maximum Likelihood) estimate by taking the derivative of the log-likelihood function, and reconstructing the number of peers $x_{i}$ as:
@@
We apply Eqn. (4) for peers (of our nodes) that 1) use Geth or Erigon (before v2.49.0), and 2) we exclude nodes with a non-public git commit string, which is a sign of software modification as we discuss in the next section. If both of our monitoring nodes observe the same peer, we take the average between two. In total, we manage to estimate the number of peer connections for 6,005 nodes: mean and median of 41 and 31, respectively. For those, each peer has sent an average (median) of 293,121 (81,327) transaction messages. To make the distribution generalized/smooth and get $g(x)$, we apply the kernel density function (KDE) based on SciPy’s implementation & Scotts’ method for the bandwidth selection. Figure 7 shows the resulting PDF (Probability Density Function) and CDF (Cumulative Density Function) for our estimate of active peer connections for each node based on data collected from our two measurement nodes. The strong alignment between the estimations between two nodes indicates our method is robust. As an additional check, we ask one Ethereum node operator to provide us with the node name, node ID, and the maximum number of peers. The node uses an unmodified Geth node and sets the maximum number of peers to 40. Our estimator tells us that this node is expected to have $33.68$ peers. Given that not all the nodes are stable over time, our estimate appears to be consistent with this piece of ground truth.
@@ 6.2 Ratio of modified nodes: $\gamma$
With regard to duplications (which are state-independent), we flag transactions as “spamming” (i.e., duplicated) transactions when we see the attacker changing the data field to produce more than 10 transactions (e.g., just tweaking 1 byte) while keeping other parameters constant. We then label peers that forward duplicated transactions within one second, which indicates that the duplicated transactions stay in each peer’s txpool at the same time. That fact suggests the absence of a gas bump rule (or that the threshold was set to zero).
@@
Based on the HEAD of the git commit in the node name (which is generated when compiling the source code), we determine whether the node uses the publicly available version of the software or the private forks of the repository. 174 out of 175 peers share an identical git commit hash, which is not present in the corresponding GitHub repository. (We fail to collect information for one remaining node.) The details of how we extract the git commit hash and use it to determine whether a node was customized can be found in Appendix E. Further, we find that these modified nodes apply an aggressive propagation strategy ($\pi$=“aggressive”) – i.e., they only broadcast (0x02), but never announce (0x08), further increasing amplification effects.
@@ 6.3 Amplification factor
To quantify the economic impact of network waste (EAF), we next calculate the financial loss of spreading invalid transactions by modified nodes. We observe that many centralized services deploy their nodes in the cloud service (Appendix F), so we calculate the data transfer cost based on the “pay-as-you-go” policy of the cloud services. We choose AWS pricing [^8] https://aws.amazon.com/ec2/pricing/on-demand/#Data_Transfer_within_the_same_AWS_Region as an example. AWS does not charge for inbound traffic but for outbound traffic (significantly less for the traffic within AWS). Most cloud services follow a similar standard. If we assume that the cost of inbound traffic is zero, the EDoS attack incurs costs only for 1) the attacker (to send invalid transactions) and 2) the modified nodes (to propagate invalid transactions to the rest).
@@ 7 Attack simulation in the local P2P network
We evaluate our EDoS attack in the local P2P network simulation and compare it with the existing DoS attacks.
@@
We set up a local P2P network, in which we deploy our Geth node (modified or unmodified), attack our own nodes, and then assess both txpool congestion and transaction inclusion in the blocks. We base our attack on Yaish et al.’s publicly available repository, [^10] https://github.com/AvivYaish/SpeculativeDoS which allows for a direct comparison of our approach with the two existing DoS attacks, namely 1) Baseline, that is, sending valid transactions with higher gas prices (naive eviction strategy), and 2) MemPurge [54], that is, sending future latent (invalid) transactions by extending the DETER method [35]. In these two attacks, we do not modify our node.
@@
We set up one validator, 80 honest accounts, and a varying number $x$ of malicious accounts, all of which interact with our node through the RPC endpoint. First, each honest account sends 64 transactions to collectively fill up the transaction pool with valid transactions. Each attacker then begins submitting attack transactions. For our EDoS attack, each attack account attempts to transfer an amount (1 ETH) higher than its current account balance, and starting from nonce 0, which is lower than the current nonce – i.e., these transactions are invalid. Each attack account then gradually increments the nonce to generate 32 invalid transactions (i.e., a total of $32x$ attack transactions) with the intention of evicting honest transactions.
@@
We measure 1) the ratio of honest transactions in the txpool, 2) the ratio of honest transactions in the final block, and 3) the number of attack transactions in the block (i.e., the attack cost). We calibrate the number of malicious accounts controlled by the attacker to measure the changes in those three metrics. Figure 10 depicts the relationship between those three metrics ($y$-axis) and the number of malicious accounts controlled by the attacker ($x$-axis).
@@
Figure 10 (left) illustrates that our EDoS attack can evict as many honest transactions as the Baseline attack. MemPurge requires even a larger number of attack accounts to replace honest transactions, as the node pushes back future transactions to the queue if the txpool is already full.
@@
Figures 10 (middle and right) illustrate that our EDoS attack can exclude honest transactions not just in the txpool but block as well because the attacker’s invalid transactions occupy the txpool. Furthermore, these invalid transactions are never included in the block, resulting in zero attack transaction costs. MemPurge gradually displaces honest transactions in the block, while the attacker pays for one transaction per attack account. The Baseline attack replaces honest transactions with a smaller number of attack accounts. Yet, all the attack transactions are included in the block, resulting in a significantly high(er) attack cost.
@@
In summary, our attack achieves the same level of txpool/block eviction rate as the two existing DoS attacks but with significantly lower (zero) attack transaction costs. This advantage allows the attacker to keep refreshing the txpool with its new invalid transactions, causing the modified node to continuously notify its neighbors, thereby incurring egress traffic loss—EDoS attack. In Appendix G, we describe more detailed attack scenarios for the EDoS attack and roughly estimate the maximum traffic monthly costs incurred at the modified node to be 88,904 USD per node (8M USD in aggregate) when an attacker saturates outgoing traffic with invalid transactions.
@@
We open-source our attack code in our repository [^11] https://github.com/taro-tsuchiya/BlockchainEDoS and document the attack implementation details.
@@ 8.1 Simulating validations
We first create an isolated fork of the Ethereum blockchain within our local network to estimate the Geth validation processing time. We use a modified Geth client, integrated with the Pebble database. [^12] Geth can run on either Pebble or LevelDB as its underlying database. This approach allows us to simulate a controlled environment without impacting the live P2P network. Our objective is to examine how different types of validations contribute to the overall processing time and identify potential areas for efficiency improvements. Specifically, we run two functions in the first “accept” stage: ValidateTransaction() and ValidateTransactionWithState() for 1,000 consecutive blocks, starting from block number 18,140,000, in a total of 136,437 transactions. Simulating the initial stage (“first”) would suffice to estimate validation time, as the node caches the results of the account fetch. We conduct our experiment on a Ubuntu 22.04.2 LTS server, equipped with an AMD Ryzen Threadripper 3990X 64-Core Processor and a Corsair MP400 NVMe disk.
@@ 9.1 Cost-benefit analysis
We summarize the benefits and the costs for each player: an attacker node ($\mathcal{A}$), a modified node ($\mathcal{M}$), and a regular node ($\mathcal{R}$). The attacker can launch an EDoS attack to induce financial losses for modified nodes. For instance, an attacker economically competing with the entity operating modified nodes could use the attack to weaken the competition. The attacker can also use its invalid transactions to displace legitimate transactions, rendering the service unusable (i.e., causing a DoS) as shown in §7.
@@
Based on §7 and §8.1, while omitting validation checks leads to a possible reduction in processing time by the order of milliseconds, the marginal time savings may not necessarily justify the potential damages/risks from a Blockchain EDoS attack (in terms of MEV profits). Modified nodes engaging in these “optimizations” also expect to face degradation in the quality of service provided, due to the increasing number of invalid transactions (as shown in §7). With this in mind, a rationally economic node would still continue skipping validations as long as the economic benefits (including other forms of payments such as user subscription fees) outweigh any financial losses incurred from the attack.

### v2

@@ 4.3 Network waste
To estimate how one invalid transaction amplifies in the network, we first derive 1) the number of bytes a regular node $i$ receives based on the number of peer connections (see blue arrows directed toward each regular node in Figure 1), 2) the number of bytes the entire network wastes based on the distributions of active connections on each peer (all the blue arrows). This calculation only considers the public nodes in a P2P network, but not the users (MEV bots and block miners/builders) who 1) privately connect to txpool providers or 2) listen to transactions through RPC endpoints outside the P2P network. The actual amplification factor would be the combination of 1) the amount of traffic that goes to users who listen to modified nodes and 2) TAF. We indeed notice that some modified nodes shut down the outbound connections to reduce outbound costs. Yet, they still incur the cost of propagating invalid transactions to users.
@@
As we already determine the waste for each regular node with a varying number of active connections, we proceed to calculate the total waste in the network by considering the distribution of peer connections, $g(x)$, which we empirically estimate in §6.1. There are $(1-\gamma)\cdot N=N_{\mathcal{R}}$ regular nodes where $N$ is the total number of nodes in the network. We multiply the amount of waste for each node ($f(x)$) and the distributions of peer connections $g(x)$ to get the overall network waste:
Network $\text{Network waste}=\sum_{i=1}^{N_{\mathcal{R}}}f(x_{i})g(x_{i})\ ,$ (2)
In our measurement, we estimate $g(x)$ for a part of the network (i.e., our connected peers), and smooth $g(x)$ into a continous form. In that case, the network waste is $N_{\mathcal{R}}\cdot\int_{0}^{1000}f(x)g(x)$ where we assume a maximum connections of 1,000 to ensure convergence and exclude an impractically large connections. We finally divide this by the transaction size ($a$) to get TAF. An adversary could set up modified nodes on their own to increase $\gamma$, but these nodes would also be attacked, making this strategy unproductive.
@@ 5 Empirical examples of attack
This section identifies the similar attack instances seen in the wild, and characterizes the size, intensity, cost, and strategies, and actors (the attacker – sender address, the victim – centralized services) involved in those attacks. While we cannot definitively pinpoint the motivations of those attackers, those attacks are executed in a manner consistent with our proposed attack (see the discussion in Appendix C).
@@ 5.1 Victims
Figure 5 (right) is a co-occurrence matrix where each cell indicates the number of dropped transactions common to both services (the $x$- and $y$-axis) normalized by each day. The diagonal (from the bottom left to the top right) is the number of dropped transactions to each source—e.g., bloXroute has around 60,000 new dropped transactions per day. bloXroute, Chainbound, and Eden collectively accept a significantly high number of dropped transactions that are not seen in any of the other txpools, which suggests that they may have employed different propagation strategies from others. Those results hypothesize that a part of the latency reduction might also come from transaction validation process, which motivates our simulations in §8.1.
Figure 5: Comparing the centralized services’ txpools (left: accuracy-latency plot, right: dropped transactions between sources.
@@ 5.3 Characterizing attack behavior
We first check the three types of invalid transactions discussed in §4.1 and Figure 3: 1) insufficient balance, 2) past nonce, and 3) duplicate transaction. To check the balance and nonce of the relevant accounts at the time of transaction submission, we use our archive node to fetch the state of the blockchain at the time of the timestamp. Given that there is a network delay between when the transaction enters a P2P network and when Flashbots records the timestamp, we look at the state of the account two blocks before the observation timestamp (i.e., at least 12 seconds in the past). Of 2,591 attack cases, we confirm that 536 instances (from 224 unique addresses) come from accounts that do not have a sufficient balance. Two services—bloXroute and Eden—have included those transactions during our observation window. In particular, some addresses do not have any transactions (send or receive) in their history whatsoever (i.e., zero Ether balance). In addition, three services (Chainbound, Eden, bloXroute) have accepted transactions that were using a past nonce (and thus, are invalid) during our measurement interval. Just like we identify invalid transactions from accounts with insufficient balance, we retrieve the last nonce of the corresponding address and identify 62 attacks from 27 unique addresses that use previously used nonce.
@@ 6.1 Peer connection distribution: $g(x)$
We formalize our method for estimating the number of active connections for each peer, $x_{i}$, and empirically computing its distribution, $g(x)$. Although studies on blockchain P2P network topology exist, the key novelty in our approach is to estimate active peer connections on the Ethereum mainnet without submitting test transactions (i.e., mediating ethical concerns and research costs). Instead, we reverse-engineer the number of peers based on propagation message types. We have 2.5 billion messages in five months (as explained in §3.2).
As outlined in §4, an unmodified Geth node $i$ has a number of peer connections $x_{i}$ and uniformly randomly chooses $\sqrt{x_{i}}$ nodes to broadcast (0x02) transactions, while sending announces (0x08) to the rest. Our node receives a 0x02 message with probability $\theta_{i}=\frac{\sqrt{x_{i}}}{x_{i}}$, which means the data distribution follows a binomial distribution with parameters $\theta_{i}$ and $m$, where $m$ is the number of messages we receive from each peer. Specifically, $m=m_{2}+m_{8}$ where $m_{2}$ and $m_{8}$ are the number of 0x02 and 0x08 messages sent by peer $i$, respectively. The probability (likelihood function) of the binomial distribution is defined as
@@
We derive an unbiased maximum likelihood estimate $\hat{\theta}_{i}$ by taking the derivative of the log-likelihood function, and reconstruct the number of peers $x_{i}$ as:
@@
We apply Eqn. (4) for peers (of our nodes) that 1) use Geth or Erigon (before v2.49.0), and 2) we exclude nodes with a non-public git commit string, which is a sign of software modification as we discuss in the next section. If both of our monitoring nodes observe the same peer, we take the average between two. In total, we manage to estimate the number of peer connections for 6,005 nodes: mean and median of 41 and 31, respectively. For those, each peer has sent an average (median) of 293,121 (81,327) transaction messages. To make the distribution generalized/smooth and get $g(x)$, we apply the kernel density function (KDE) based on SciPy’s implementation & Scotts’ method for the bandwidth selection. Figure 7 shows the resulting PDF (Probability Density Function) and CDF (Cumulative Density Function) of active peer connections (green: estimate from Node A, orange: from Node B). The strong alignment between the estimations between two nodes indicates our method is robust. As an additional check, we ask one Ethereum node operator to provide us with the node name, node ID, and the maximum number of peers. The node uses an unmodified Geth node and sets the maximum number of peers to 40. Our estimator tells us that this node is expected to have $33.68$ peers. Given that not all the nodes are stable over time, our estimate appears to be consistent with this piece of ground truth.
@@ 6.2 Ratio of modified nodes: $\gamma$
With regard to duplications, we prepare the set of duplicated transactions (e.g., changing the data field while keeping other parameters constant.) We then label peers that forward two duplicated transactions within one second, which indicates that the duplicated transactions stay in each peer’s txpool at the same time. That fact suggests the absence of a gas bump rule (or that the threshold was set to zero).
@@
Based on the HEAD of the git commit in the node name (which is generated when compiling the source code), we determine whether the node uses the publicly available version of the software or the private forks of the repository. 174 out of 175 peers share an identical git commit hash, which is not present in the corresponding GitHub repository. (We fail to collect information for one remaining node.) This fact ensures that an invalid transaction reaches out to all modified nodes. The details of how we extract the git commit hash and use it to determine whether a node was customized can be found in Appendix E. Further, we find that these modified nodes apply an aggressive propagation strategy ($\pi$=“aggressive”) – i.e., they only broadcast (0x02), but never announce (0x08), further increasing amplification effects.
@@ 6.3 Amplification factor
To quantify the economic impact of network waste (EAF), we next calculate the financial loss of spreading invalid transactions by modified nodes. We observe that many centralized services deploy their nodes in the cloud service (Appendix F), so we calculate the data transfer cost based on the “pay-as-you-go” policy of the cloud services. We choose AWS pricing [^8] https://aws.amazon.com/ec2/pricing/on-demand/#Data_Transfer_within_the_same_AWS_Region as an example. AWS does not charge for inbound traffic but for outbound traffic (plus, significantly less for the traffic within AWS). Most cloud services follow a similar standard. If we assume that the cost of inbound traffic is zero, the attack incurs costs only for 1) the attacker (to send invalid transactions) and 2) the modified nodes (to propagate invalid transactions to the rest).
@@ 7 Attack simulation in the local P2P network
We evaluate our proposed attack in the local P2P network simulation and compare it with the existing DoS attacks.
@@
We set up a local P2P network, in which we deploy our modified node, attack the node, and then assess both txpool congestion and transaction inclusion in the blocks. We base our attack on Yaish et al.’s publicly available repository, [^10] https://github.com/AvivYaish/SpeculativeDoS which allows for a direct comparison of our approach with the two existing DoS attacks, namely 1) Baseline, that is, sending valid transactions with higher gas prices (naive eviction strategy), and 2) MemPurge [55], that is, sending future latent (invalid) transactions by extending the DETER method [36]. We test all three attacks in our modified node.
@@
We set up one validator, 80 honest accounts, and a varying number $x$ of malicious accounts, all of which interact with our node through the RPC endpoint. First, each honest account sends 64 transactions to collectively fill up the transaction pool with valid transactions. Each attacker then begins submitting attack transactions. For our proposed attack, each attack account attempts to transfer an amount higher than its current account balance, and starting from nonce 0, which is lower than the current nonce – i.e., these transactions are invalid. Each attack account then gradually increments the nonce to generate 32 invalid transactions (i.e., a total of $32x$ attack transactions) with the intention of evicting honest transactions.
@@
We measure 1) the ratio of honest transactions in the txpool, 2) the ratio of honest transactions in the final block, and 3) the number of attack transactions in the block (i.e., the attack transaction cost). We calibrate the number of malicious accounts controlled by the attacker to measure the changes in those three metrics. Figure 10 depicts the relationship between those three metrics ($y$-axis) and the number of malicious accounts controlled by the attacker ($x$-axis).
@@
Figure 10 (left) illustrates that our proposed attack can evict as many honest transactions as the Baseline attack. MemPurge requires even a larger number of attack accounts to replace honest transactions, as the node pushes back future transactions to the queue if the txpool is already full.
@@
Figures 10 (middle and right) show that our proposed attack can exclude honest transactions not just in the txpool but block as well because the attacker’s invalid transactions occupy the txpool. Furthermore, these invalid transactions are never included in the block, resulting in zero attack transaction costs. MemPurge gradually displaces honest transactions in the block, while its attacker pays for one transaction per attack account. The Baseline attack replaces honest transactions with a smaller number of attack accounts. Yet, all the attack transactions are included in the block, resulting in a significantly high(er) attack cost.
@@
In summary, our attack achieves the same level of txpool/block eviction rate as the two existing DoS attacks but with significantly lower (zero) attack transaction costs. This advantage allows the attacker to keep refreshing the txpool with its new invalid transactions, causing the modified node to continuously notify its neighbors, thereby incurring egress traffic loss—EDoS attack. In Appendix G, we describe more detailed attack scenarios and roughly estimate the maximum traffic monthly costs incurred at the modified node to be 88,904 USD per node (8M USD in aggregate) when an attacker saturates outgoing traffic with invalid transactions.
@@
We open-source our attack code in our repository111111hhttps://github.com/taro-tsuchiya/BlockchainAmplification and document the attack implementation details.
@@ 8.1 Simulating validations
We first create an isolated fork of the Ethereum blockchain within our local network to estimate the Geth validation processing time. We use a modified Geth client, integrated with the Pebble database. [^12] Geth can run on either Pebble or LevelDB as its underlying database. This approach allows us to simulate a controlled environment without impacting the live P2P network. Our objective is to examine how different types of validations contribute to the overall processing time and identify potential areas for efficiency improvements. Specifically, we run two functions in the first “accept” stage: ValidateTransaction() and ValidateTransactionWithState() for 1,000 consecutive blocks, starting from block number 18,140,000, in a total of 136,437 transactions. We conduct our experiment on a Ubuntu 22.04.2 LTS server, equipped with an AMD Ryzen Threadripper 3990X 64-Core Processor and a Corsair MP400 NVMe disk.
@@ 9.1 Cost-benefit analysis
We summarize the benefits and the costs for each player: an attacker, modified nodes, and regular nodes. The attacker can launch an attack to induce financial losses for modified nodes. For instance, an attacker economically competing with the entity operating modified nodes could use the attack to weaken the competition. The attacker can also use its invalid transactions to displace legitimate transactions, rendering the service unusable (i.e., causing a DoS) as shown in §7.
@@
Based on §7 and §8.1, while omitting validation checks leads to a possible reduction in processing time by the order of milliseconds, the marginal time savings may not necessarily justify the potential damages/risks from the Blockchain Amplification Attack (in terms of MEV profits). Modified nodes engaging in these “optimizations” also expect to face degradation in the quality of service provided, due to the increasing number of invalid transactions (as shown in §7). With this in mind, a rationally economic node would still continue skipping validations as long as the economic benefits (including revenue from users paying to connect to those services) outweigh any financial losses incurred from the attack.